OpenStudy (dschneider2016):

Please help!!! Find Medal!!! Find the cube roots of 8(cos 216° + i sin 216°)

OpenStudy (dschneider2016):

@jdoe0001 @whpalmer4

OpenStudy (dschneider2016):

@satellite73

OpenStudy (anonymous):

what is the cubed root of 8?

OpenStudy (dschneider2016):

2

OpenStudy (anonymous):

good, what is \(216\div3\)?

OpenStudy (dschneider2016):

72

OpenStudy (anonymous):

then one answer is \[2\left(\cos(72^\circ)+i\sin(72^\circ)\right)\] easy right?

OpenStudy (dschneider2016):

yes! how do I find the other ones?

OpenStudy (anonymous):

ready to find the other ones?

OpenStudy (anonymous):

what is \(216+360\)? (we are going around the circle again)

OpenStudy (dschneider2016):

576

OpenStudy (anonymous):

ok i believe you what is \(576\div3\)?

OpenStudy (dschneider2016):

192

OpenStudy (anonymous):

ok so answer 2 is \[2\left(\cos(192^\circ)+i\sin(192^\circ)\right)\]

OpenStudy (anonymous):

lather, rinse, repeat you only have to do it one more time

OpenStudy (dschneider2016):

great! So would I add 720 to 216 and then divide by 3 to find the next one

OpenStudy (anonymous):

yes

OpenStudy (dschneider2016):

2(cos(312∘)+isin(312∘))

OpenStudy (anonymous):

same as dividing the unit circle in to three parts, with 72 as one of them, but it is just as easy to keep adding 360

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

see how easy that is?

OpenStudy (dschneider2016):

Great! SO is that all?

OpenStudy (dschneider2016):

Thank you so much!!

OpenStudy (anonymous):

yes that is pretty much the whole point of writing \(a+bi\) in trig form otherwise this would be very very hard

OpenStudy (anonymous):

yw